HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

MENLO SYSTEMS’ SOLUTIONS FOR HIGH PRECISION SPECTROSCOPY

\underline{A. AUBOURG}, S. DE VEGA, Menlo Systems GmbH, Christina Mariel, Bunsenstr. 5, 82152 Martinsried, Germany

Menlo Systems is a leading developer and global supplier of instrumentation for precision metrology. The company, with headquarters in Martinsried near Munich, is known for its Nobel Prize winning optical frequency comb technology. Our main product lines are optical frequency combs, solutions for time and frequency distribution, ultrastable lasers, terahertz systems, and femtosecond lasers.

Recently, Menlo Systems introduced two new products for applications in the important fingerprint region for high accuracy spectroscopy or nanoFTIR. The “YLMO Mid-I” is a femtosecond fiber-laser with a tunable wavelength over 3–5 μm and the “Mid-IR Comb” provides stabilized modes (enabling dual-comb spectroscopy setups) from 3 μm up to 14 μm.

p-number: p260  Submitted on Wed, 05 Jul 2023 16:57:41 +0200